A proton travels in a circle perpendicular to a magnetic field. The field strength is 2T and the proton speed is 20m/s. Calculate the following:

radius of circle

period of motion

frequency of motion
Answer #1

Radius of circle r=\frac{mv}{eB}=\frac{1.6726*10^{-27}*20}{1.6*10^{-19}*2}=1.045*10^{-7}\,m

Period of motion T=\frac{2\pi r}{v}=\frac{2\pi *1.045*10^{-7}}{20}=3.28*10^{-8}\,s

Frequency of motion f=\frac{1}{T}=\frac{1}{3.28*10^{-8}}=3.04*10^7\,Hz